Determine the solution set for the following equations.
7x+2z=21
-51-y-3z=5
4x+4y-z=4

Answer by ikleyn(52867)

.
Determine the solution set for the following equations.
7x+2z=21
-51-y-3z=5
4x+4y-z=4
~~~~~~~~~~~~~~~~


Double check your 2-nd equation.

It looks strange and suspicious.


            Well-wisher